The nuclei of large atoms, such as uranium, with 92protons, can be modeled as spherically symmetric spheres of charge. The radius of the uranium nucleus is approximately 7.4×10−15m .

a)What magnitude of electric field does it produce at the distance of the electrons, which is about 1.9×10−10 m ?

b)The electrons can be modeled as forming a uniform shell of negative charge. What net electric field do they produce at the location of the nucleus?
Answer #1

Charge on nucleus (Q) = 92x1.6x10-19=1.472x10-17 C
radius of atom (R) = 7.4x10-15m

a) Electric field at the distance of the electrons

E = kQ/r2

using the value of k = 9 x 109 Nm2/C2 , r = 1.9 x 10-10 m

E = 6.9726x 1012 N/C
Hence magnitude of electric field does it produce at the distance of the electrons(r = 1.9 x 10-10 m), is E = 6.9726x 1012 N/C

b) Net electric field produced by electrons at the location of the nucleus = zero

(Because inside a shell of charge ,the electric field is equal to zer